# A Comparison of Relativistic Coupled Cluster and Equation of Motion Coupled Cluster Quadratic Response Theory

### v1

@@ Abstract
We present the implementation of relativistic coupled cluster quadratic response theory (QR-CC), following our development of relativistic equation of motion coupled cluster quadratic response theory (QR-EOMCC) [X. Yuan et al., J. Chem. Theory Comput. 2023, 19, 9248]. These codes, which can be used in combination with relativistic (2- and 4-component based) as well as non-relativistic Hamiltonians, are capable of treating both static and dynamic perturbations for electric and magnetic operators. We have employed this new implementation to revisit the calculation of static and frequency-dependent first hyperpolarizabilities of hydrogen halides (HX, X=F-Ts) and the Verdet constant of heavy noble gas atoms (Xe, Rn, Og), in order to investigate the differences and similarities of QR-CC and the more approximate QR-EOMCC. Furthermore, we have determined the relative importance of scalar relativistic effects and spin-orbit coupling to these properties, through a comparison of different Hamiltonians, and extended our calculations to superheavy element species (HTs for hyperpolarizabilities, Og for the Verdet constant). Our results show that as one moves towards the bottom of the periodic table, QR-EOMCC can yield rather different results (hyperpolarizabilities) or perform rather similarly (Verdet constant) to QR-CC. These results underscore the importance of further characterizing the performance of QR-EOMCC for heavy element systems.
@@ 1 Introduction
External perturbations such as those induced by electric or magnetic fields allow us to probe the electronic structure of atoms, molecules and materials, with different perturbations providing us with complementary information about the system 1, 2. The first-order (linear) responses to such a perturbation are particularly sought-after quantities, since from them one can obtain information about electronically or vibrationally 3 excited states though processes involving a single photon, or magnetic properties such as NMR shieldings or indirect spin-spin coupling constants 4, 5, 6 to name but a few.
@@
Properties arising from the interaction with two or more photons 7, 8 at a time are also of great interest be it for technological applications in non-linear optics 9, 10 or for speciation through two-photon absorption 11, as the differences in selection rules compared to one-photon processes allow for probing different states.
@@
Theoretical models are indispensable to interpret experimental results, and the last fifty years have seen the development 12, 13 of approaches based on ab initio molecular electronic structure theory gain popularity in applications, thanks in no small part to the development of response theory 14, 15, 2 and its application to different models such as multi-configurational self-consistent field theory 16, 17, 18, density functional theory 19, and coupled cluster theory 20, 21, 22. The use of the time-averaged quasienergy formalism 23, 24, 25, 21, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35 has provided not only a unified framework, with which to derive response for both variational and non-variational approaches, but also a way to enable the development of embedding approaches in which variational and non-variational electronic structure methods are combined 36, 37, 38, 39, 40, 41, 42, 43 to calculate molecular properties in large systems within a fully quantum mechanical description in a cost-effective manner.
@@
In the case of coupled cluster theory, one can identify two main approaches, one which employs an exponential parametrization for the time evolution of the cluster operators (CC response), and another based on the equation of motion formalism (EOMCC response)44, 45, 46, 47, 48, 49, 50, 51, 52, 53 for which a linear parametrization is employed to obtain molecular properties.
@@
From the outset, key differences between the two approaches have been identified with respect to separability 21, 54, 55, with as consequence that CC transition moments are size-extensive whereas EOMCC ones are not. More recently the interconnections between these two approaches were further investigated from a formal point of view 56, 57.
@@
In spite of such differences, the calculation of response properties within the EOMCC formalism remains appealing from both the implementation and the application points of view 58, 59, 60, as simplifications in the equations also translate into a somewhat more modest computational cost. This calls for an understanding of how the two approaches differ in practical calculations. In the literature comparisons between linear response CC (LR-CC) and EOMCC (LR-EOMCC) properties 61, 49, 50, 57 so far suggest that, in spite of the lack of size-extensivity of EOM-CC transition moments, in practice the two approaches yield quite similar results.
@@
Similar conclusions have been reached when comparing quadratic response CC (QR-CC) and EOMCC (QR-EOMCC) results for dipole hyperpolarizabilties 48, 46, 57, though comparisons were initially limited to a few systems. More recently, an extensive benchmark study on excited state transition moments 62, which also probe quadratic response functions, seem to confirm the tendencies observed previously. These previous evaluations were carried out for valence properties of molecules containing light elements, that is, occupying the first two rows of the periodic table. For properly representing the molecular properties of heavier systems it is necessary to account for relativistic effects63, 64, 65, 66, 67, 68, 69 and to the best of our knowledge there is currently no data in the literature assessing the differences between the two approaches beyond the second row of the periodic table.
@@
Response theory can be used in combination with relativistic Hamiltonians70, and implementations for mean-field approaches (Hartree-Fock and DFT) are for example available for the calculation of frequency-dependent linear and non-linear response properties for 4-component-based Hamiltonians71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81. To go beyond mean-field, relativistic coupled cluster theory 82, 83, 84, 85, 86, 87 offers a flexible framework for obtaining accurate energies and properties which treats electron correlation and relativistic effects on the same footing. It can thereby be used to investigate systems of interest across the periodic table, and benchmark more approximate methods.
@@
In recent years there have been extensive efforts by different groups to address states beyond closed-shell ground states through EOMCC theory for excitation and ionization energies, and electron affinities88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 103, 107, 108, 109, 110, 111, 112. While some of these implementations allow for the calculation of transition moments or static properties (through the application of finite fields or via approximate expectation values), full-fledged implementations of CC or EOMCC response theory are much more scarce. Yuan et al. 113 have reported an implementation capable of carrying out frequency-dependent LR-CC and LR-EOMCC calculations for different (from 1- to 4-component based) Hamiltonians for electric, magnetic and mixed electric-magnetic properties. A comparison of CC and EOMCC showed that for dipole polarizabilities the two approaches do yield similar results as relativistic effects become more pronounced, but for indirect spin-spin coupling constants there is growing difference between the two. Chakraborty et al. 114 reported an efficient implementation of LR-CC for the spin-free eXact 2-Component Hamiltonian (x2C) and applied it to investigate the static and dynamic dipole polarizabilities of an extensive set of molecular systems.
@@
Beyond linear response, the analytical treatment of quadratic response coupled cluster theory would be particularly useful in applications in which finite field calculations are not straightforward, easy to converge or inapplicable, such for frequency-dependent calculations of non-linear optical properties or in the study of birefringences. However, general implementations beyond linear response are scarcer still, with our report of relativistic QR-EOMCC 115 being, to the best of our knowledge, the first to appear in the literature. This implementation allowed us to gauge the performance of more approximate approaches such as TDDFT and TDHF in the calculation of hyperpolarizabilies, two-photon absorption cross-sections and Verdet constants. However, as discussed above QR-CC and QR-EOMCC are not equivalent, and it is desirable to further characterize the behavior of the latter with respect to the former across the periodic table, before carrying out more extensive applications with the more approximate QR-EOMCC scheme.
@@
The aim of the current work is therefore to report our implementation of quadratic response coupled cluster for relativistic Hamiltonians, capable of handling for both static and frequency-dependent perturbations involving electric or magnetic operators. We showcase it by providing a first comparison of how QR-CC and QR-EOMCC behave as one goes towards the bottom of the periodic table. We thereby chose to re-investigate the dipole first hyperpolarizabilities of hydrogen halides, and the Verdet constants of heavy noble gas atoms as they show significant relativistic effects 115. Going further than this previous work, we carried out a more extensive exploration on the effect of the Hamiltonian on these properties, and investigated the properties of systems made up of superheavy elements, such as the Tennessine hydride molecule and the Oganesson atom.
@@
Tennessine and Oganesson, whose existence has been recognized by IUPAC and IUPAC in 2015, are the two heaviest elements known. They have been synthesized at the Joint Institute for Nuclear Research in Dubna by nuclear fusion upon bombardment of Berkelium and Californium targets by between atoms of Calcium-48 ions, respectively116, 117. While they are expected to be members of groups 17 and 18, the difficulties in their production–only a few atoms are generated at a time, and their lifetimes are in the order of milliseconds–present formidable challenges to the experimental characterization of their physico-chemical properties 118, making theoretical work an indispensable tool to learn about their electronic structure 69. In view of their large polarizabilties and more pronounced spin-orbit coupling than that in 6p, 5d and 5f block elements, they are also of potential interest as probes for any subtle difference between CC and EOMCC in the calculation of response properties.
@@ 2.1 Quadratic Response Theory
As was the case in our previous work on coupled cluster linear response113 (LR-CC) and EOM-CC quadratic response 115 (QR-EOMCC), our work is carried out within the framework of the quasienergy formulation of response theory 31. For the sake of brevity, we only review the formalism for CC quadratic response (QR-CC), and refer to the supplementary information for the working equations of QR-EOMCC as presented in 115. Throughout the text, the CC acronym will imply the use of coupled cluster singles and doubles (CCSD) wavefunctions.
We shall remain within the no-pair approximation 83, 119, 120, with the electronic part of the unperturbed molecular Hamiltonian given by
where and denote general (occupied or positive energy virtual) indexes, denotes matrix elements of the one-body operator , and the anti-symmetrized matrix elements of the two-body operator . In the calculations reported in this work, corresponds to Coulomb operator. In our current implementation, Gaunt or Breit interactions are assumed to be treated in an effective manner121, 122, 123, 124. External time-(in)dependent perturbations will be represented by one-body operators
@@
Due to the fact that in CC response theory the time-dependent quasienergy Lagrangian 31 is made stationary with respect to the wavefunction parametrization (CC amplitudes and multipliers), we can use the () and () rules in perturbation theory31, 13. The calculation of a third-order property such as quadratic response, thus only requires the determination of the unperturbed and first-order perturbed CC amplitudes and Lagrange multipliers.
@@
with for particle lines and for hole lines. As a shorthand notation, in the following we shall use greek letters to represent the strings of second quantization operators, with subscripts to represent single or double excitation; for instance, whereas will represent a generic (either single or double) excitation operator. Conversely, will denote a de-excitation operator, which is biorthogonal to operator , that is, , with representing the reference state, here a single Slater determinant made up of molecular orbitals–or molecular spinors whenever the Hamiltonian accounts for spin-orbit coupling. Also, denotes the ensemble of excited determinants with respect to the reference.
From the exponential ansatz for the CC wavefunction 125, 126
@@
where is matrix representation of the normal-ordered similarity transformed Hamiltonian. The definitions of and are given in table 1. Further details on the evaluation of these equations in our implementation can be found in 84.
After solving for the unperturbed wave function parameters, the first-order parameters are next obtained by solving the response equations 31, 113, 115
where is the identity matrix. We employ the shorthand notation to describe the first-order perturbed amplitudes and multipliers, obtained for perturbation at frequency , respectively. We note that, in contrast to the first-order response equations in QR-EOMCC, the equations for the first-order perturbed multipliers depend on the first-order perturbed amplitudes. The definitions of terms , are also given in table 1.
The frequency-(in)dependent quadratic response function is given by the expression
where the perturbing frequencies satisfy the relationship . The operator takes care of the permutation of the perturbing operators, and
is used for the symmetrization of the response functions 31, which requires the solution of response also for the negative of the frequencies (inverting the frequencies followed by complex conjugation). Eq. 11 can also be rewritten as
where and group the terms involving contractions only with perturbed amplitudes or with both perturbed amplitudes and multipliers, respectively:
and that for all combinations allowed by the product.
As is the case for and 113, 115, the , , and tensors, given by the expression in the table 1, are never constructed explicitly but rather as contractions with the perturbed amplitudes and/or multipliers. The programmable expressions required to assemble the term are presented in the supplementary information. The programmable expressions for , and can be obtained from minor modifications of those already presented in our previous work 113, 115.

### v3

@@ Abstract
We present the implementation of relativistic coupled cluster quadratic response theory (QR-CC), following our development of relativistic equation of motion coupled cluster quadratic response theory (QR-EOMCC) [X. Yuan et al., J. Chem. Theory Comput. 2023, 19, 9248–9259]. These codes, which can be used in combination with relativistic (2- and 4-component based) as well as non-relativistic Hamiltonians, are capable of treating both static and dynamic perturbations for electric and magnetic operators. We have employed this new implementation to revisit the calculation of static and frequency-dependent first hyperpolarizabilities of hydrogen halides (HX, X=F-Ts) and the Verdet constant of heavy noble gas atoms (Xe, Rn, Og) and of selected hydrogen halides (HF to HI), in order to investigate the differences and similarities of QR-CC and the more approximate QR-EOMCC. Furthermore, we have determined the relative importance of scalar relativistic effects and spin-orbit coupling to these properties, through a comparison of different Hamiltonians, and extended our calculations to superheavy element species (HTs for hyperpolarizabilities, Og for the Verdet constant). Our results show that as one moves towards the bottom of the periodic table, QR-EOMCC can yield rather different results (hyperpolarizabilities) or perform rather similarly (Verdet constant) to QR-CC. These results underscore the importance of further characterizing the performance of QR-EOMCC for heavy element systems.
@@ 1 Introduction
External perturbations such as those induced by electric or magnetic fields allow us to probe the electronic structure of atoms, molecules and materials, with different perturbations providing us with complementary information about the system norman_principles_2018, Jaszuski2015. The first-order (linear) responses to such a perturbation are particularly sought-after quantities, since from them one can obtain information about electronically or vibrationally bishop_molecular_1990 excited states though processes involving a single photon, or magnetic properties such as NMR shieldings or indirect spin-spin coupling constants helgaker1999ab, vaara2007theory, helgaker2008quantum to name but a few.
@@
Properties arising from the interaction with two or more photons papadopoulos_non-linear_2006, he2008multiphoton at a time are also of great interest be it for technological applications in non-linear optics Norman2006, boyd2020nonlinear or for speciation through two-photon absorption cronstrand_multi-photon_2005, as the differences in selection rules compared to one-photon processes allow for probing different states.
@@
Theoretical models are indispensable to interpret experimental results, and the last fifty years have seen the development helgaker_recent_2012, helgaker_molecular_2014 of approaches based on ab initio molecular electronic structure theory gain popularity in applications, thanks in no small part to the development of response theory norman_perspective_2011, Pedersen2015, Jaszuski2015 and its application to different models such as multi-configurational self-consistent field theory dalgaard1980time, olsen_linear_1985, Hettema1992, density functional theory Casida1996, and coupled cluster theory Monkhorst2009, dalgaard_aspects_1983, Koch1990a, Koch1990b, Koch1994, kobayashi1994calculation, kondo_orthogonally_1995, piecuch_property_1995. The use of the time-averaged quasienergy formalism langhoff1972aspects, rice1991calculation, sasagane1993higher, Koch1994, Christiansen1995a, Christiansen1995b, Christiansen1996, Hattig1997, coriani_coupled_1997, Christiansen1998, Gauss1998, kauczor2013communication, khani_uv_2019, hattig_cc2_2000 has provided not only a unified framework, with which to derive response for both variational and non-variational approaches, but also a way to enable the development of embedding approaches in which variational and non-variational electronic structure methods are combined Hofener2012, Hofener2012b, Hofener2013, Hofener2016, Niemeyer2025, Myhre2014, Myhre2016, Goletto2021 to calculate molecular properties in large systems within a fully quantum mechanical description in a cost-effective manner.
@@
In the case of coupled cluster theory, one can identify two main approaches, one which employs an exponential parametrization for the time evolution of the cluster operators (CC response)Monkhorst2009, dalgaard_aspects_1983, Koch1990a, Koch1990b, Koch1994, kobayashi1994calculation, Christiansen1995a, Christiansen1995b, Christiansen1996, Hattig1997, coriani_coupled_1997, Christiansen1998, kondo_orthogonally_1995, piecuch_property_1995, spirko_molecular_1996, piecuch_molecular_1996, piecuch_convergence_1997, and another based on the equation of motion formalism (EOMCC response)Stanton1993, Stanton1993b, Sekino1995, gauss_coupledcluster_1995, Rozyczko1997, Sekino1999, Crawford2009, Nanda2015, nanda_static_2016, nanda_communication_2018 for which a linear parametrization is employed to obtain molecular properties.
@@
From the outset, key differences between the two approaches have been identified with respect to separability Koch1994, Stanton1994, caricato2009difference, with as consequence that CC transition moments are size-intensive whereas EOMCC ones are not. More recently the interconnections between these two approaches were further investigated from a formal point of view pawlowski_molecular_2015, coriani_molecular_2016.
@@
In spite of such differences, the calculation of response properties within the EOMCC formalism remains appealing from both the implementation and the application points of view faber2018resonant, andersen2022cherry, andersen2022probing, as simplifications in the equations also translate into a somewhat more modest computational cost. This calls for an understanding of how the two approaches differ in practical calculations. In the literature comparisons between linear response CC (LR-CC) and EOMCC (LR-EOMCC) properties Sekino1994, Sekino1999, Crawford2009, coriani_molecular_2016 so far suggest that, in spite of the lack of size-extensivity of EOM-CC transition moments, in practice the two approaches yield quite similar results.
@@
Similar conclusions have been reached when comparing quadratic response CC (QR-CC) and EOMCC (QR-EOMCC) results for dipole hyperpolarizabilties Rozyczko1997, Sekino1995, coriani_molecular_2016, though comparisons were initially limited to a few systems. More recently, an extensive benchmark study on excited state transition moments irek2025, which also probe quadratic response functions, seem to confirm the tendencies observed previously. These previous evaluations were carried out for valence properties of molecules containing light elements, that is, occupying the first two rows of the periodic table. For properly representing the molecular properties of heavier systems it is necessary to account for relativistic effectsPyykko1988a, saue_relativistic_2011, Pyykk2012, Autschbach2012a, liu2017handbook, Schwerdtfeger2020, Smits2023 and to the best of our knowledge there is currently no data in the literature assessing the differences between the two approaches beyond the second row of the periodic table.
@@
Response theory can be used in combination with relativistic HamiltoniansSaue2002_PostDirac, and implementations for mean-field approaches (Hartree-Fock and DFT) are for example available for the calculation of frequency-dependent linear and non-linear response properties for 4-component-based Hamiltoniansvisscher_4-component_1997, Saue:Jensen:JCP2003, norman_quadratic_2004, salek2005linear, Henriksson:2005, Henriksson_JCP2006, Tellgren_JCP2007, bast:JCP2009, Bast_ijqc2009, Bast2009b, Villaume_JCP2010. To go beyond mean-field, relativistic coupled cluster theory Liu2021, visscher1996formulation, pototschnig_implementation_2021, Yuwono2024, Chamoli2025, Brandejs2025 offers a flexible framework for obtaining accurate energies and properties which treats electron correlation and relativistic effects on the same footing. It can thereby be used to investigate systems of interest across the periodic table, and benchmark more approximate methods.
@@
In recent years there have been extensive efforts by different groups to address states beyond closed-shell ground states through EOMCC theory for excitation and ionization energies, and electron affinitiesPathak:2016ck, Pathak:2015he, Pathak:2014gv, Pathak:2014dm, Pathak:2016dk, Klein:2008hx, Yang:2012gd, Wang:2015jw, Epifanovsky:2015hsa, Cao:2016fx, Cao:2017id, Zhang:2017jj, Akinaga:2017hv, Wang:2016hx, shee2018equation, Koulias2019, Halbert2021, Halbert2023, Zhang2024, Koulias2019, Zhang2024b, Uhlov2024, Yuwono2025, Li2025, Banerjee2025, Mukhopadhyay2025. While some of these implementations allow for the calculation of transition moments or static properties (through the application of finite fields or via approximate expectation values), full-fledged implementations of CC or EOMCC response theory are much more scarce. Yuan_Xiang_LR have reported an implementation capable of carrying out frequency-dependent LR-CC and LR-EOMCC calculations for different (from 1- to 4-component based) Hamiltonians for electric, magnetic and mixed electric-magnetic properties. A comparison of CC and EOMCC showed that for dipole polarizabilities the two approaches do yield similar results as relativistic effects become more pronounced, but for indirect spin-spin coupling constants there is growing difference between the two. Chakraborty2025 reported an efficient implementation of LR-CC for the spin-free eXact 2-Component Hamiltonian (X2C) and applied it to investigate the static and dynamic dipole polarizabilities of an extensive set of molecular systems.
@@
Beyond linear response, the analytical treatment of quadratic response coupled cluster theory would be particularly useful in applications in which finite field calculations are not straightforward, easy to converge or inapplicable, such for frequency-dependent calculations of non-linear optical properties or in the study of birefringences. However, general implementations beyond linear response are scarcer still, with our report of relativistic QR-EOMCC Yuan_Xiang_QR being, to the best of our knowledge, the first to appear in the literature. This implementation allowed us to gauge the performance of more approximate approaches such as TDDFT and TDHF in the calculation of hyperpolarizabilies, two-photon absorption cross-sections and Verdet constants. However, as discussed above QR-CC and QR-EOMCC are not equivalent, and it is desirable to further characterize the behavior of the latter with respect to the former across the periodic table, before carrying out more extensive applications with the more approximate QR-EOMCC scheme.
@@
The aim of the current work is therefore to report our implementation of quadratic response coupled cluster for relativistic Hamiltonians, capable of handling for both static and frequency-dependent perturbations involving electric or magnetic operators. We showcase it by providing a first comparison of how QR-CC and QR-EOMCC behave as one goes towards the bottom of the periodic table. We thereby chose to re-investigate the first dipole hyperpolarizabilities of hydrogen halides, and the Verdet constants of heavy noble gas atoms as they show significant relativistic effects Yuan_Xiang_QR and include an exploration of the Verdet constants of selected hydrogen halides. Going further than this previous work, we carried out a more extensive exploration on the effect of the Hamiltonian on these properties, and investigated the properties of systems made up of superheavy elements, such as the Tennessine hydride molecule and the Oganesson atom.
@@
Tennessine and Oganesson, whose existence has been recognized by IUPAC and IUPAP in 2015, are the two heaviest elements known. They have been synthesized at the Joint Institute for Nuclear Research in Dubna by nuclear fusion upon bombardment of Berkelium and Californium targets by between atoms of Calcium-48 ions, respectivelyKarol2016a, Karol2016b. While they are expected to be members of groups 17 and 18, the difficulties in their production–only a few atoms are generated at a time, and their lifetimes are in the order of milliseconds–present formidable challenges to the experimental characterization of their physico-chemical properties Smits2023b, making theoretical work an indispensable tool to learn about their electronic structure Smits2023. In view of their large polarizabilties and more pronounced spin-orbit coupling than that in 6p, 5d and 5f block elements, they are also of potential interest as probes for any subtle difference between CC and EOMCC in the calculation of response properties.
@@ 2.1 Quadratic Response Theory
As was the case in our previous work on coupled cluster linear responseYuan_Xiang_LR (LR-CC) and EOM-CC quadratic response Yuan_Xiang_QR (QR-EOMCC), our work is carried out within the framework of the quasienergy formulation of response theory Christiansen1998. For the sake of brevity, we only review the formalism for CC quadratic response (QR-CC), and refer to the SI for the working equations of QR-EOMCC as presented in Yuan_Xiang_QR. Throughout the text, the CC acronym will imply the use of coupled cluster singles and doubles (CCSD) wavefunctions.
We shall remain within the no-pair approximation visscher1996formulation, liu_going_2013, Almoukhalalati2016, with the electronic part of the unperturbed molecular Hamiltonian given by
where and denote general (occupied or positive energy virtual) indexes, denotes matrix elements of the one-body operator , and the anti-symmetrized matrix elements of the two-body operator . In the calculations reported in this work, corresponds to Coulomb operator. In our current implementation, Gaunt or Breit interactions are assumed to be treated in an effective mannerSchimmelpfennig.Gropen.1998f8, sikkema2009molecular, iliavs2007infinite, saue2002four. External time-(in)dependent perturbations will be represented by one-body operators
@@
Due to the fact that in CC response theory the time-dependent quasienergy Lagrangian Christiansen1998 is made stationary with respect to the wavefunction parametrization (CC amplitudes and multipliers), we can use the () and () rules in perturbation theoryChristiansen1998, helgaker_molecular_2014. The calculation of a third-order property such as quadratic response, thus only requires the determination of the unperturbed and first-order perturbed CC amplitudes and Lagrange multipliers.
@@
with for particle lines and for hole lines. As a shorthand notation, in the following we shall use greek letters to represent the strings of second quantization operators, with subscripts to represent single or double excitation; for instance, whereas will represent a generic (either single or double) excitation operator. Conversely, will denote a de-excitation operator, which is biorthogonal to operator , that is, , with representing the reference state, here a single Slater determinant made up of molecular orbitals–or molecular spinors whenever the Hamiltonian accounts for spin-orbit coupling. Also, will denote an excited determinant () contained within the singles and doubles manifold.
From the exponential ansatz for the CC wavefunction shavitt2009many, bartlett2007coupled
@@
where is matrix representation of the normal-ordered similarity transformed Hamiltonian. The definitions of and are given in Table 1. Further details on the evaluation of these equations in our implementation can be found in pototschnig_implementation_2021.
After solving for the unperturbed wave function parameters, the first-order parameters are next obtained by solving the response equations Christiansen1998, Yuan_Xiang_LR, Yuan_Xiang_QR
Iyt^_, y¯t^_^Yω_YF,ξ^Y, η^Yω_Z = -(ω_X + ω_Y)P( X(ω_X), Y(ω_Y), Z(ω_Z))f^XYZ(ω_X,ω_Y,ω_Z)O^X,YO^¯X,YC^±ω P ¯HFF^X,A^Y,GBG xt^_ yt^_ zt^_F^X,A^YB